Scores on a standard test are normally distributed with a mean of 38.7 and a standard deviation of 7. Find the 90th percentile of score.

please show your works with explanation
Answer #1

Given that,

mean = = 38.7

standard deviation = =7

Using standard normal table,

P(Z < z) = 90%

= P(Z < z) = 0.90  

= P(Z < 1.28) = 0.90

z = 1.28 Using standard normal table,

Using z-score formula  

x= z * +

x= 1.28*7+38.7

x= 47.66

x=48